OpenStudy (anonymous):

Five years ago a house was worth $210,000. Now the house is worth $260,000. Assume a linear relationship between time (measured in years) and value (measured in dollars). Find a formula for the value, V(t), at time t. (t = 0 refers to now.)

OpenStudy (anonymous):

use the equation of a line is y=m*x+b. Use the value of the house as the y value and t as the x value. b represents the value of the house at t=0. for example, at t=0, 260,000=m*(0)+b, This means that b= 260,000. now plug in 210,000=m*(-5)+260,000 to solve for m. Solve this equation and m=10,000. This means that the value of the house increases by 10,000 every year.